OpenStudy (anonymous):

complete the table and write an equation that represents it. X Y -1 1 0 3 1 5 2 7 3 4

OpenStudy (solomonzelman):

you are missing something in your table (the 2 y values) .

OpenStudy (anonymous):

you are supposed to fill them in.

OpenStudy (freckles):

notice the 2nd column are what kinds of numbers?

OpenStudy (solomonzelman):

ohh.

OpenStudy (anonymous):

i'm not sure @freckles . im not good at these problems. i have no clue what im doing lol

OpenStudy (freckles):

do you know what odd numbers are?

OpenStudy (anonymous):

yes.

OpenStudy (freckles):

do you see that pattern for the second column?

OpenStudy (anonymous):

so the next two would be 9 and 11?

OpenStudy (freckles):

yep yep :)

OpenStudy (anonymous):

okay and then im not sure how to do the equation

OpenStudy (freckles):

anyways this is a linear equation find the slope and the y-intercept to find the slope take any of your two points and do \[m=\frac{y_2-y_1}{x_2-x_1}\] and then to find the y-intercept find what y is when x=0 (your tabled actually gives you this value)<--we will call this b so once you find m and b plug into y=mx+b

OpenStudy (freckles):

so one thing at a time let's find the slope

OpenStudy (freckles):

which two points do you want to use it will not matter will will get the same slope no matter which two you choose

OpenStudy (anonymous):

3 and 9

OpenStudy (freckles):

(3,9) is one point

OpenStudy (freckles):

we need one more more

OpenStudy (anonymous):

1 and 5

OpenStudy (freckles):

ok cool (3,9) and (1,5) are the two points you have chose and we will use

OpenStudy (freckles):

(3,9) -(1,5) ----- <-subtract these points

OpenStudy (freckles):

9-5=? 3-1=?

OpenStudy (anonymous):

4, 2

OpenStudy (freckles):

(3,9)-(1,5)=(2,4) (2,4) will give us the slope if we do y/x=4/2=?

OpenStudy (anonymous):

okay. im not sure how to do that.

OpenStudy (freckles):

4/2?

OpenStudy (freckles):

4/2=what 2*what=4

OpenStudy (anonymous):

2

OpenStudy (freckles):

yes you do know how to divide :)

OpenStudy (freckles):

y=mx+b m is the slope and we just found it to be 2 so we have y=2x+b

OpenStudy (freckles):

we still need to find b

OpenStudy (freckles):

this is easy though

OpenStudy (freckles):

because to find b (the y-intercept) you just find what y is when x=0

OpenStudy (freckles):

look at your table

OpenStudy (anonymous):

3?

OpenStudy (freckles):

look at the first column until you see x=0 and then go on that row to find what the y is that corresponds to that x=0 yes when x=0 we have y=3 3 is therefore b

OpenStudy (freckles):

so you have found your equation for this line y=2x+3

OpenStudy (anonymous):

so that is my eqation?

OpenStudy (anonymous):

equation*

OpenStudy (freckles):

yes

OpenStudy (anonymous):

ok. thank you!

OpenStudy (freckles):

np do you have any questions about this particular problem?

OpenStudy (freckles):

or on anything we did here?

OpenStudy (anonymous):

no i understand now.

OpenStudy (freckles):

cool